# Error-Correcting Codes for Two Bursts of $t_{1}$-Deletion-$t_{2}$-Insertion with Low Computational Complexity

### v1

@@ Abstract
Burst errors involving simultaneous insertions, deletions, and substitutions occur in practical scenarios, including DNA data storage and document synchronization, motivating developments of channel codes that can correct such errors. In this paper, we address the problem of constructing error-correcting codes (ECCs) capable of handling multiple bursts of $t_{1}$-deletion-$t_{2}$-insertion ($(t_{1},t_{2})$-DI) errors, where each burst consists of $t_{1}$ deletions followed by $t_{2}$ insertions in a binary sequence. We make three key contributions: Firstly, we establish the fundamental equivalence of (1) two bursts of $(t_{1},t_{2})$-DI ECCs, (2) two bursts of $(t_{2},t_{1})$-DI ECCs, and (3) one burst each of $(t_{1},t_{2})$-DI and $(t_{2},t_{1})$-DI ECCs. Then, we derive lower and upper bounds on the code size of two bursts of $(t_{1},t_{2})$-DI ECCs, which can naturally be extended to the case of multiple bursts. Finally, we present constructions of two bursts of $(t_{1},t_{2})$-DI ECCs. Compared to the codes obtained by the syndrome compression technique, the resulting codes achieve significantly lower computational complexity.
@@ I Introduction
Despite many recent works, the development of ECCs for synchronization errors is a fundamental challenge. Going back to the foundations, the seminal work by Levenshtein [11] constructed a single-edit (insertion/deletion/substitution) correcting code requiring only $\log n+1$ bits of redundancy, using the well-known Varshamov-Tenengolts (VT) syndrome [33]. [^1] In this paper, unless stated otherwise, all logarithmic operations are base two. While there were only sporadic results for a long time, recent breakthroughs by Brakensiek et al. [3] on $t$-deletion codes have spurred substantial progress in this field [6, 25, 7, 23, 21, 22, 17, 30, 4, 28, 13, 29, 19] over the last decade. For the specific case of two deletions, the state-of-the-art result was achieved by Guruswami et al. [7] for binary codes with $4\log n+o(\log n)$ bits of redundancy. This result was extended to non-binary alphabets by Song et al. [28] and Liu et al. [13]. They introduced codes with $5\log n+o(\log n)$ bits of redundancy under certain parameter constraints. For general $t$-deletions with $t\geq 2$, Sima et al. [21, 23] developed binary codes with $4t\log n+o(\log n)$ bits of redundancy using the syndrome compression technique [24]. Subsequent work by Song et al. [29] improved this result to $(4t-1)\log n+o(\log n)$ through pre-coding methods. Regarding the $q$-ary case, Sima et al. [22] initially achieved at most $30t\log n+o(\log n)$ bits of redundancy for $q$-ary $t$-deletion codes, and Schaller et al. [19] recently improved this result to $t\log q+(3t-1)\log n+o(\log n)$ bits of redundancy for sufficiently large $n$ and $q>n^{2+\epsilon}$, where $\epsilon>0$.
@@
In some practical scenarios, data corruption may occur as bursts of insertions, deletions, and substitutions. Motivated by this, significant research has focused on designing ECCs capable of handling such errors as well. Several works have investigated code constructions for correcting a single burst of $t$-deletions or up to $t$-deletions, including [20, 10, 28, 27, 34]. Among these, the recent coding scheme proposed by Song et al. [27] achieves a redundancy of $\log n+8\log\log n+o(\log\log n)$ bits for correcting a single burst of at most $t$-deletions. As a further extension in this direction, Sun et al. [31] proposed ECCs with $\log n+o(\log n)$ bits of redundancy to correct a single burst of edits in $q$-ary sequences ($q\geq 2$). Beyond the single-burst setting, Sima et al. [24] constructed ECCs capable of correcting $m$ bursts of at most $t$-deletions with a redundancy of $4m\log n+o(\log n)$ bits by employing the syndrome compression technique. After that, focusing on the case of two bursts of exactly $t$-deletions, Ye et al. [35] presented explicit ECC constructions with $5\log n+o(\log n)$ bits of redundancy.
@@
In real-world applications, multiple error types may occur simultaneously. Early work by Smagloy et al. [26] addressed this challenge by constructing binary ECCs correcting a single deletion and a single substitution. Later, Song et al. [29] generalized this result to multiple deletions and multiple substitutions under both binary and non-binary settings. A notable contribution by Schoeny et al. [20] introduced burst ECCs capable of handling both deletions and insertions simultaneously, specifically constructing binary codes for one burst of 22-deletions-11-insertion (denoted by $(2,1)$-DI) errors. This result was later extended by Lu et al. to one burst of $(t_{1},1)$-DI ECCs [14] and subsequently to $(t_{1},t_{2})$-DI ECCs with $t_{1}\geq 2t_{2}$ [15], thereby establishing a lower bound of $\log(n-t_{1}+2)+t_{1}-1$ bits of redundancy for such codes. Most recently, Sun et al. [32] provided explicit constructions of $q$-ary one burst of $(t_{1},t_{2})$-DI ECCs with $\log n+O(1)$ bits of redundancy, applicable for any non-negative integers $t_{1},t_{2}$ and $q\geq 2$.
@@
Building upon the developments for correcting different types of burst errors, in this paper, we first establish the equivalence of two bursts of $(t_{1},t_{2})$-DI ECCs, two bursts of $(t_{2},t_{1})$-DI ECCs, and one burst of $(t_{1},t_{2})$-DI one burst of $(t_{2},t_{1})$-DI ECCs. We then derive both lower and upper bounds on the cardinality of two bursts of $(t_{1},t_{2})$-DI ECCs. Based on these results, we further propose novel constructions of binary ECCs capable of correcting two bursts of $(t_{1},t_{2})$-DI with $t_{1},t_{2}$ being two positive integers. The resulting codes achieve substantially reduced computational complexity compared to those obtained by the syndrome compression technique.
@@ III-B The Bounds on the Cardinality of Two Bursts of $(t_{1},t_{2})$-DI ECCs
$a_{1}=2^{2t_{2}-2}\left({n-2t_{1}+2\choose 2}+1\right),$
$a_{2}=2^{2t_{1}-2}\left({n-2t_{2}+2\choose 2}+1\right).$
@@
$\leq$ $2^{n-2t_{1}+2t_{2}}.$
Denoting $a_{1}=2^{2t_{2}-2}\left({n-2t_{1}+2\choose 2}+1\right)$ and $a_{2}=2^{2t_{1}-2}\left({n-2t_{2}+2\choose 2}+1\right)$, the equations above imply that
$|\mathcal{C}|\leq\frac{2^{n-2t_{1}+2t_{2}}}{\max(a_{1},a_{2})},$
@@
For any ECC $\mathcal{C}\in\Sigma_{2}^{n}$ capable of correcting two bursts of $(t_{1},t_{2})$-DI, the redundancy of $\mathcal{C}$ satisfies $2\log n+O(1)\leq r(\mathcal{C})\leq 4\log n+o(\log n)$.
@@
The upper and lower bounds on the code size of two bursts of $(t_{1},t_{2})$-DI ECCs can be naturally extended to the case of $m$ bursts with $m\geq 2$, yielding the redundancy bounds of a code $\mathcal{C}\in\Sigma_{2}^{n}$ capable of correcting $m$ bursts of $(t_{1},t_{2})$-DI errors $m\log n+O(1)\leq r(\mathcal{C})\leq 2m\log n+o(\log n)$.
@@ IV Two bursts of $(t_{1},t_{2})$-DI ECCs by Syndrome Compression Techniques
when $t_{1},t_{2}\geq 2$, and
$\begin{aligned} \left|\mathcal{N}^{DI}_{2,(t_{1},t_{2})}(\mathbf{x})\right|=&{n-2t_{1}+2\choose 2}{n-2t_{1}+2t_{2}+1\choose 2}\cdot 2^{2t_{1}-4}\\ \leq&n^{4}2^{2t_{1}-4}\end{aligned}$
when $t_{1}>t_{2}=1$.
@@
where $t\geq 1$ is a constant with respect to $n$. The cardinality of $\mathcal{N}^{DS}_{2,(1,t)}$ is
@@
V Two Bursts of $(t_{1},t_{2})$-DI ECCs
@@ V-A Two Bursts of $(t,t)$-DI ECCs
Denote $\tilde{\mathbf{x}}$ is a $2^{t-1}$-ary vector corresponding to $X_{t-1}$. Specifically, let
@@
Let $\mathbf{x}\in\Sigma_{2}^{n}$ be a binary sequence of length $n$ and $\tilde{\mathbf{x}}$ be the corresponding $2^{t-1}$-ary vector defined in (24), where $t\geq 2$ is a positive integer. If a $(t,t)$-DI occurs in $\mathbf{x}$, then there is a $(2,2)$-DI in $\tilde{\mathbf{x}}$.
@@
For any $\mathbf{x}\in\Sigma_{2}^{n}$ and an integer $t\geq 2$, two bursts of $(t,t)$-DI in $\mathbf{x}$ can be corrected by an $n/(t-1),n/(t-1)-8,d_{\mathrm{min}}=9]_{q}$ Generalized Reed-Solomon (GRS) code with at most $8\log n+O(1)$ bits of redundancy, where $q\geq n$.
@@
According to Lemma 4, it is clear that two bursts of $(t,t)$-DI in the original sequence $\mathbf{x}$ induce two bursts of $(2,2)$-DI in the transformed sequence $\tilde{\mathbf{x}}$. These errors manifest as two bursts of two symbol substitutions in $\tilde{\mathbf{x}}$, which together amount to a total of four symbol substitutions. Such error patterns can be corrected by using a GRS code with the minimum Hamming distance $9$.
Since $\tilde{\mathbf{x}}\in\Sigma_{2^{t-1}}^{n/(t-1)}$, the GRS code must be defined over a finite field $\mathbb{F}_{q}$ with $q\geq\max\big(2^{t-1},n/(t-1)\big)$. As $t\geq 2$ being a constant independent of $n$, it suffices to choose $q$ as the smallest prime satisfying $q\geq n$. Hence, an $n/(t-1),n/(t-1)-8,d_{\mathrm{min}}=9]_{q}$ GRS code with $q\geq n$ can be directly applied. As illustrated in [18], a GRS code over an alphabet of size $q\geq n$ for correcting $t$-symbol errors requires $2t\log q$ bits of redundancy. Consequently, correcting four substitution errors incurs a redundancy with at most $8\log n+O(1)$ bits. ∎
@@
When $t=1$, two bursts of $(1,1)$-DI in $\mathbf{x}\in\Sigma_{2}^{n}$ cause two substitutions in $\mathbf{x}$, then applying an $n,n-4,d_{\mathrm{min}}=5]_{q}$ GRS code with at most $4\log n+O(1)$ bits of redundancy is sufficient, where $q\geq n$.
@@ V-B Two Bursts of $(t_{1},t_{2})$-DI ECCs
In this subsection, we consider the case of $t_{1}>t_{2}$ and address design of ECCs correcting two bursts of $(t_{1},t_{2})$-DI errors. By representing a sequence in an matrix formula, we can demonstrate that two bursts of $(t_{1},t_{2})$-DI cause two bursts of $(1,t^{\prime}-1)$-DS in each row of the matrix, where $t^{\prime}=\lceil t_{1}/(t_{1}-t_{2})\rceil$. Then, by locating the positions of the two bursts in the first row, the error positions of the original sequence can be approximately identified. Finally, two bursts of $(t_{1},t_{2})$-DI in the original sequence can be corrected following the function proposed in Construction 1 with reduced computational complexity.
@@
Meanwhile, if $j_{1}-i_{2}-t+1$ is even, then we have
@@
    One can find two intervals $\mathcal{J}_{1},\mathcal{J}_{2}\subseteq[n]$ with length $|\mathcal{J}_{1}|,|\mathcal{J}_{2}|\leq(d\log(n/(t_{1}-t_{2}))+2t^{\prime})(t_{1}-t_{2})$ such that one burst of $(t_{1},t_{2})$-DI occurs in $\mathbf{x}_{\mathcal{J}_{1}}$ and the other in $\mathbf{x}_{\mathcal{J}_{2}}$, respectively.
@@
    One can find an interval $\mathcal{J}\subseteq[n]$ with length $|\mathcal{J}|\leq(3d\log{(n/(t_{1}-t_{2}))}+4t^{\prime})(t_{1}-t_{2})$ such that two bursts of $(t_{1},t_{2})$-DI occur in $\mathbf{x}_{\mathcal{J}}$.
@@
$\mathcal{J}_{1}=\left[(i-t^{\prime})(t_{1}-t_{2})+2,(i^{\prime}+t^{\prime}-1)(t_{1}-t_{2})\right],$
$\mathcal{J}_{2}=\left[(j-t^{\prime})(t_{1}-t_{2})+2,(j^{\prime}+t^{\prime}-1)(t_{1}-t_{2})\right].$
@@
$|\mathcal{J}_{1}|,|\mathcal{J}_{2}|\leq(d\log{n\over t_{1}-t_{2}}+2t^{\prime})(t_{1}-t_{2}).$
@@
$\mathcal{J}=\left[(i-t^{\prime})(t_{1}-t_{2})+2,(i^{\prime}+t^{\prime}-1)(t_{1}-t_{2})\right].$
@@
$|\mathcal{J}|\leq\left(3d\log{n\over t_{1}-t_{2}}+4t^{\prime}\right)(t_{1}-t_{2}).$
@@
For any binary sequence $\mathbf{x}\in\Sigma_{2}^{n}$ and two positive integers $t_{1},t_{2}$ with $t_{1}>t_{2}$, denote by $X_{t_{1}-t_{2}}$ the matrix representation of $\mathbf{x}$ following (25), and $X[1]$ the first row of $X_{t_{1}-t_{2}}$. Let $f_{1}(\mathbf{x}),f_{2}(\mathbf{x})$ and $f(\mathbf{x})$ be the functions capable of correcting one burst of $(t_{1},t_{2})$-DI, two bursts of $(t_{1},t_{2})$-DI, and two bursts of $(1,t^{\prime}-1)$-DS proposed in Construction 1 and Lemma 3 for any $\mathbf{x}\in\Sigma_{2}^{n}$, respectively. Let $\mathcal{L}_{j}$ and $\mathcal{L}_{j^{\prime}}$ be the intervals defined in (35) with $\rho_{1}=(d\log(n/(t_{1}-t_{2}))+2t^{\prime})(t_{1}-t_{2})$ and $\rho_{2}=(3d\log(n/(t_{1}-t_{2}))+4t^{\prime})(t_{1}-t_{2})$, respectively, where $j\in[\lceil n/\rho_{1}\rceil-1/],j^{\prime}\in[\lceil n/\rho_{2}\rceil-1]$. For each $\mathbf{x}\in\Sigma_{2}^{n}$ with $X[1]$ being a $d$-regular sequence, $a\in[0,2]$ and $i\in[0,1]$, denote
@@
Case 1. Lemma 7-a) holds. In this case, we can find two intervals $\mathcal{J}_{1}$ and $\mathcal{J}_{2}$ with $|\mathcal{J}_{1}|,|\mathcal{J}_{2}|\leq\rho_{1}=(d\log(n/(t_{1}-t_{2}))+2t^{\prime})(t_{1}-t_{2})$ such that two bursts occur in $\mathbf{x}_{\mathcal{J}_{1}}$ and $\mathbf{x}_{\mathcal{J}_{2}}$, respectively. Following (35), there exist two intervals $\mathcal{L}_{j_{1}},\mathcal{L}_{j_{2}}$ such that $\mathcal{J}_{1}\subseteq\mathcal{L}_{j_{1}},\mathcal{J}_{2}\subseteq\mathcal{L}_{j_{2}}$, where $j_{1},j_{2}\in[\lceil n/\rho_{1}\rceil-1]$. Without loss of generality, let $j_{1}\leq j_{2}$ in the following.

### v2

@@ Abstract
Burst errors involving simultaneous insertions, deletions, and substitutions occur in practical scenarios, including DNA data storage and document synchronization, motivating developments of channel codes that can correct such errors. In this paper, we construct error-correcting codes (ECCs) capable of handling multiple bursts of $t_{1}$-deletion-$t_{2}$-insertion ($(t_{1},t_{2})$-DI) errors, where each burst consists of $t_{1}$ deletions followed by $t_{2}$ insertions in a binary sequence. We make three key contributions: Firstly, we establish the fundamental equivalence of (1) two bursts of $(t_{1},t_{2})$-DI ECCs, (2) two bursts of $(t_{2},t_{1})$-DI ECCs, and (3) one burst each of $(t_{1},t_{2})$-DI and $(t_{2},t_{1})$-DI ECCs. Then, we derive lower and upper bounds on the code size of two bursts of $(t_{1},t_{2})$-DI ECCs, which can naturally be extended to the case of multiple bursts. Finally, we present constructions of two bursts of $(t_{1},t_{2})$-DI ECCs. Compared to the codes obtained by the direct application of the syndrome compression technique, the resulting codes achieve significantly lower computational complexity.
@@ I Introduction
Despite many recent developments, the design of ECCs for synchronization errors remains as a fundamental challenge. Going back to the foundations, the seminal work by Levenshtein [11] constructed a single-edit (insertion/deletion/substitution) correcting code requiring only $\log n+1$ bits of redundancy, using the well-known Varshamov-Tenengolts (VT) syndrome [32]. [^1] In this paper, unless stated otherwise, all logarithmic operations are base two. While there were only sporadic results for a long time, recent breakthroughs by Brakensiek et al. [3] on $t$-deletion codes have spurred substantial progress in this field [6, 24, 7, 22, 20, 21, 17, 29, 4, 27, 13, 28, 18] over the last decade. For the specific case of two deletions, the state-of-the-art result was achieved by Guruswami et al. [7] for binary codes with $4\log n+o(\log n)$ bits of redundancy. This result was extended to non-binary alphabets by Song et al. [27] and Liu et al. [13], which introduced codes with $5\log n+o(\log n)$ bits of redundancy under certain parameter settings. For the general case of $t$-deletions with $t\geq 2$, Sima et al. [20, 22] developed binary codes with $4t\log n+o(\log n)$ bits of redundancy using the syndrome compression technique [23]. Subsequent work by Song et al. [28] improved this result to $(4t-1)\log n+o(\log n)$ through pre-coding methods. Regarding the $q$-ary case, Sima et al. [21] initially achieved a redundancy of at most $30t\log n+o(\log n)$ bits for $q$-ary $t$-deletion codes, and Schaller et al. [18] recently improved this result to $t\log q+(3t-1)\log n+o(\log n)$ bits of redundancy for sufficiently large $n$ and $q>n^{2+\epsilon}$, where $\epsilon>0$.
@@
In some practical scenarios, data corruption may occur as bursts of insertions, deletions, and substitutions. Motivated by this, significant research has focused on designing ECCs capable of handling such errors as well. Several works have investigated code constructions for correcting a single burst of $t$-deletions or up to $t$-deletions, including [19, 10, 27, 26, 33]. Among them, the recent coding scheme proposed by Song et al. [26] achieves a redundancy of $\log n+8\log\log n+o(\log\log n)$ bits for correcting a single burst of at most $t$-deletions. As a further extension in this direction, Sun et al. [30] proposed ECCs with $\log n+o(\log n)$ bits of redundancy to correct a single burst of edits in $q$-ary sequences ($q\geq 2$). Beyond the single-burst setting, Sima et al. [23] constructed ECCs capable of correcting $m$ bursts of at most $t$-deletions with a redundancy of $4m\log n+o(\log n)$ bits by employing the syndrome compression technique. Subsequently, focusing on the case of two bursts of exactly $t$-deletions, Ye et al. [34] presented explicit ECC constructions with $5\log n+o(\log n)$ bits of redundancy.
@@
In real-world applications, multiple error types may occur simultaneously. Early work by Smagloy et al. [25] addressed this challenge by constructing binary ECCs correcting a single deletion and a single substitution. Later, Song et al. [28] generalized this result to multiple deletions and multiple substitutions under both binary and non-binary settings. A notable contribution by Schoeny et al. [19] introduced burst ECCs capable of handling both deletions and insertions simultaneously, specifically constructing binary codes for one burst of 22-deletions-11-insertion (referred to as $(2,1)$-DI) errors. This result was later extended by Lu et al. to one burst of $(t_{1},1)$-DI ECCs [14] and subsequently to $(t_{1},t_{2})$-DI ECCs with $t_{1}\geq 2t_{2}$ [15], thereby establishing a lower bound of $\log(n-t_{1}+2)+t_{1}-1$ bits of redundancy for such codes. Most recently, Sun et al. [31] provided explicit constructions of $q$-ary one burst of $(t_{1},t_{2})$-DI ECCs with $\log n+O(1)$ bits of redundancy, applicable for any non-negative integers $t_{1},t_{2}$ and $q\geq 2$.
@@
Building upon the developments for correcting different types of burst errors, in this paper, we first establish the equivalence of two bursts of $(t_{1},t_{2})$-DI ECCs, two bursts of $(t_{2},t_{1})$-DI ECCs, and one burst of $(t_{1},t_{2})$-DI one burst of $(t_{2},t_{1})$-DI ECCs. We then derive both lower and upper bounds on the cardinality of two bursts of $(t_{1},t_{2})$-DI ECCs. Based on these results, we propose novel constructions of binary ECCs capable of correcting two bursts of $(t_{1},t_{2})$-DI with $t_{1},t_{2}$ being two positive integers. The resulting codes achieve substantially reduced computational complexity compared to those obtained by the direction application of the syndrome compression technique.
@@ III-B The Bounds on the Cardinality of Two Bursts of $(t_{1},t_{2})$-DI ECCs
$a_{1}=2^{2t_{1}-2}\left({n-2t_{1}+2\choose 2}+1\right),$
$a_{2}=2^{2t_{2}-2}\left({n-2t_{2}+2\choose 2}+1\right).$
@@
$\leq$ $2^{n-2t_{2}+2t_{1}}.$
Denoting $a_{1}=2^{2t_{1}-2}\left({n-2t_{1}+2\choose 2}+1\right)$ and $a_{2}=2^{2t_{2}-2}\left({n-2t_{2}+2\choose 2}+1\right)$, the equations above imply that
$|\mathcal{C}|\leq\frac{2^{n}}{\max(a_{1},a_{2})},$
@@
For any ECC $\mathcal{C}\subseteq\Sigma_{2}^{n}$ capable of correcting two bursts of $(t_{1},t_{2})$-DI, the redundancy of $\mathcal{C}$ satisfies $2\log n+O(1)\leq r(\mathcal{C})\leq 4\log n+o(\log n)$.
@@
The upper and lower bounds on the code size of two bursts of $(t_{1},t_{2})$-DI ECCs can be naturally extended to the case of $m$ bursts with $m\geq 2$, yielding the redundancy bounds of a code $\mathcal{C}\subseteq\Sigma_{2}^{n}$ capable of correcting $m$ bursts of $(t_{1},t_{2})$-DI errors $m\log n+O(1)\leq r(\mathcal{C})\leq 2m\log n+o(\log n)$.
@@ IV Two bursts of $(t_{1},t_{2})$-DI ECCs by Syndrome Compression Techniques
when $t_{1},t_{2}\geq 2$, and when $t_{1}>t_{2}=1$,
$\begin{aligned} \left|\mathcal{N}^{DI}_{2,(t_{1},t_{2})}(\mathbf{x})\right|=&{n-2t_{1}+2\choose 2}{n-2t_{1}+2t_{2}+1\choose 2}\cdot 2^{2t_{1}-4}\\ \leq&n^{4}2^{2t_{1}-4}.\end{aligned}$
@@
where $t\geq 1$. The cardinality of $\mathcal{N}^{DS}_{2,(1,t)}$ is
@@
V Two Bursts of $(t_{1},t_{2})$-DI ECCs with Low Complexity
@@ V-A Two Bursts of $(t,t)$-DI ECCs
Denote a $2^{t-1}$-ary vector corresponding to $X_{t-1}$ by
@@
Let $\mathbf{x}$ be a binary sequence of length $n$ and $\tilde{\mathbf{x}}$ be the corresponding $2^{t-1}$-ary vector defined in (24), where $t\geq 2$ is a positive integer. If a burst of $(t,t)$-DI occurs in $\mathbf{x}$, then there is a burst of $(2,2)$-DI in $\tilde{\mathbf{x}}$.
@@
For any $\mathbf{x}\in\Sigma_{2}^{n}$ and an integer $t\geq 2$, define
$f^{\mathbf{u}}(\mathbf{x})=\{\mathbf{x}\in\Sigma_{2}^{n}:\tilde{\mathbf{x}}\in\{\mathbf{y}+\mathbf{u}:\mathbf{y}\in\mathcal{C}_{1}\},$
where $\mathcal{C}_{1}$ is an $n/(t-1),n/(t-1)-8,d_{\mathrm{min}}=9]_{q}$ Reed-Solomon (RS) code with $q\geq n/(t-1)$ being a prime, and $\mathbf{u}\in\mathbb{F}_{q}^{n/(t-1)}$. Then, two bursts of $(t,t)$-DI in $\mathbf{x}$ can be corrected by $f^{\mathbf{u}}(\mathbf{x})$ and $\mathbf{x}^{\prime}\in\mathcal{B}_{2,(t,t)}^{DI}(\mathbf{x})$ with at most $8\log n+O(1)$ bits of redundancy.
@@
According to Lemma 4, it is clear that two bursts of $(t,t)$-DI in the original sequence $\mathbf{x}$ induce two bursts of $(2,2)$-DI in the transformed sequence $\tilde{\mathbf{x}}$. These errors manifest as two bursts of two symbol substitutions in $\tilde{\mathbf{x}}$, which together amount to a total of four symbol substitutions. Such error patterns can be corrected since $\tilde{\mathbf{x}}\in\{\mathbf{y}+\mathbf{u}:\mathbf{y}\in\mathcal{C}_{1}\}$ and $d_{\mathrm{min}}(\mathcal{C}_{1})=9$.
Since $\mathcal{C}_{1}$ is a subspace of $\mathbb{F}_{q}^{n}$ with dimension n8n-8, its $q^{8}$ cosets form a partition of $\mathbb{F}^{n}_{q}$. That is, there exist vectors $\mathbf{u}_{i}\in\mathbb{F}_{q}^{n/(t-1)},i\in[q^{8}]$ such that $\mathcal{C}_{1}^{i}\triangleq\{\mathbf{y}+\mathbf{u}_{i}:\mathbf{y}\in\mathcal{C}_{1}\}$, and $\mathcal{C}_{1}^{i},i\in[q^{8}]$ constitutes a partition of $\mathbb{F}_{q}^{n}$. Note that each coset $\mathcal{C}_{1}^{i}$ has the same error-correcting capabilities as $\mathcal{C}_{1}$ since $d_{\mathrm{min}}(\mathcal{C}_{1}^{i})=d_{\mathrm{min}}(\mathcal{C}_{1})$. Thus, for each coset $\mathcal{C}_{1}^{i}$, a corresponding function $f^{\mathbf{u}_{i}}(\mathbf{x})$ can be constructed. By the pigeonhole principle, there exists an $i\in[q^{8}]$ such that
$|f^{\mathbf{u}_{i}}(\mathbf{x})|\geq\frac{\bigcup_{i\in[q^{8}]}|f^{\mathbf{u}_{i}}(\mathbf{x})|}{q^{8}}.$
Consequently, there exists a choice of $\mathbf{u}_{i}$ for which the corresponding coset contributes at most $8\log n+O(1)$ bits of redundancy, which is also the redundancy of $f^{\mathbf{u}}(\mathbf{x})$. ∎
@@
When $t=1$, two bursts of $(1,1)$-DI in $\mathbf{x}\in\Sigma_{2}^{n}$ cause two substitutions in $\mathbf{x}$, then applying a coset of an $n,n-4,d_{\mathrm{min}}=5]_{q}$ RS code with a prime $q\geq n$ is sufficient, which results in at most $4\log n+O(1)$ bits of redundancy.
@@ V-B Two Bursts of $(t_{1},t_{2})$-DI ECCs
In this subsection, we consider the case of $t_{1}>t_{2}$ and construct ECCs capable of correcting two bursts of $(t_{1},t_{2})$-DI errors. By representing a sequence in an matrix formula, we can demonstrate that two bursts of $(t_{1},t_{2})$-DI cause two bursts of $(1,t^{\prime}-1)$-DS in each row of the matrix, where $t^{\prime}=\lceil t_{1}/(t_{1}-t_{2})\rceil$. Then, by locating the positions of the two bursts in the first row, the error positions of the original sequence can be approximately identified. Finally, two bursts of $(t_{1},t_{2})$-DI in the original sequence can be corrected following the function proposed in Construction 1 with reduced computational complexity.
@@
Meanwhile, if $j_{1}-i_{2}-t+1$ is even, we have
@@
    One can find two intervals $\mathcal{J}_{1},\mathcal{J}_{2}\subseteq[n]$ with length $|\mathcal{J}_{1}|,|\mathcal{J}_{2}|\leq(d\log(n/(t_{1}-t_{2}))+2t^{\prime}+1)(t_{1}-t_{2})$ such that one burst of $(t_{1},t_{2})$-DI occurs in $\mathbf{x}_{\mathcal{J}_{1}}$ and the other in $\mathbf{x}_{\mathcal{J}_{2}}$, respectively.
@@
    One can find an interval $\mathcal{J}\subseteq[n]$ with length $|\mathcal{J}|\leq(3d\log{(n/(t_{1}-t_{2}))}+4t^{\prime}+1)(t_{1}-t_{2})$ such that two bursts of $(t_{1},t_{2})$-DI occur in $\mathbf{x}_{\mathcal{J}}$.
@@
$\mathcal{J}_{1}=\left[(i-t^{\prime}-1)(t_{1}-t_{2})+2,(i^{\prime}+t^{\prime}-1)(t_{1}-t_{2})\right],$
$\mathcal{J}_{2}=\left[(j-t^{\prime}-1)(t_{1}-t_{2})+2,(j^{\prime}+t^{\prime}-1)(t_{1}-t_{2})\right].$
@@
$|\mathcal{J}_{1}|,|\mathcal{J}_{2}|\leq(d\log{n\over t_{1}-t_{2}}+2t^{\prime}+1)(t_{1}-t_{2}).$
@@
$\mathcal{J}=\left[(i-t^{\prime}-1)(t_{1}-t_{2})+2,(i^{\prime}+t^{\prime}-1)(t_{1}-t_{2})\right].$
@@
$|\mathcal{J}|\leq\left(3d\log{n\over t_{1}-t_{2}}+4t^{\prime}+1\right)(t_{1}-t_{2}).$
@@
For any $\mathbf{x}\in\Sigma_{2}^{n}$ and two positive integers $t_{1},t_{2}$ with $t_{1}>t_{2}$, denote by $X_{t_{1}-t_{2}}$ the matrix representation of $\mathbf{x}$ following (25), and $X[1]$ the first row of $X_{t_{1}-t_{2}}$. Let $f_{1}(\mathbf{x}),f_{2}(\mathbf{x})$ and $f(\mathbf{x})$ be the functions capable of correcting one burst of $(t_{1},t_{2})$-DI, two bursts of $(t_{1},t_{2})$-DI, and two bursts of $(1,t^{\prime}-1)$-DS proposed in Construction 1 and Lemma 3 for any $\mathbf{x}\in\Sigma_{2}^{n}$, respectively. Let $\mathcal{L}_{j}$ and $\mathcal{L}_{j^{\prime}}$ be the intervals defined in (35) with $\rho_{1}=(d\log(n/(t_{1}-t_{2}))+2t^{\prime}+1)(t_{1}-t_{2})$ and $\rho_{2}=(3d\log(n/(t_{1}-t_{2}))+4t^{\prime}+1)(t_{1}-t_{2})$, respectively, where $j\in[\lceil n/\rho_{1}\rceil-1/],j^{\prime}\in[\lceil n/\rho_{2}\rceil-1]$. For each $\mathbf{x}\in\Sigma_{2}^{n}$ with $X[1]$ being a $d$-regular sequence, $a\in[0,2]$ and $i\in[0,1]$, denote
@@
Case 1. Lemma 7-a) holds. In this case, we can find two intervals $\mathcal{J}_{1}$ and $\mathcal{J}_{2}$ with $|\mathcal{J}_{1}|,|\mathcal{J}_{2}|\leq\rho_{1}=(d\log(n/(t_{1}-t_{2}))+2t^{\prime}+1)(t_{1}-t_{2})$ such that two bursts occur in $\mathbf{x}_{\mathcal{J}_{1}}$ and $\mathbf{x}_{\mathcal{J}_{2}}$, respectively. Following (35), there exist two intervals $\mathcal{L}_{j_{1}},\mathcal{L}_{j_{2}}$ such that $\mathcal{J}_{1}\subseteq\mathcal{L}_{j_{1}},\mathcal{J}_{2}\subseteq\mathcal{L}_{j_{2}}$, where $j_{1},j_{2}\in[\lceil n/\rho_{1}\rceil-1]$. Without loss of generality, let $j_{1}\leq j_{2}$ in the following.